Xolisile and Xolani are twins. To reward them for passing matric well, their parents gave them 

R20 000,00 each. Each one of them decided to invest three quarters of their money for a period of 

three years. Xolisile invested her money at a simple interest rate of 6,5% per annum. Xolani invested 

her money at an interest rate of 6,5% per annum, compounded annually. What is the difference 

between the values of the two investments at the end of the period?


1
Expert's answer
2021-08-02T16:53:00-0400

"3/4 \space of\space 20000=R15000"

for Xolisile

Interest, "I=\frac{PRT}{100}=\frac{15000\times 6.5\times 3}{100}=2925"

Amount at the end of 3yrs=15000+2925=R17925

for Xolani

"A=P(1+r/100)^n=15000(1+6.5/100)^3=15000\times 1.065^3=R18119.24"

Difference at the end of 3years =18119.24-17925=R194.24